Help me find the Area of a cirlce with a shaded spot. The outer ring of the circle is the shaded spot. From center of cirlce to outside ring of circle is 2.3 cm, from center to inside of shaded circle is 1.9cm.
Answer by stanbon(75887)

Help me find the Area of a circle with a shaded spot. The outer ring of the circle is the shaded spot. From center of circle to outside ring of circle is 2.3 cm, from center to inside of shaded circle is 1.9cm.
-------------------------
Draw the picture of one circle inside the other.
Larger circle area = (pi)2.3^2
Inner circle area = (pi)1.9^2
---------------------------------------
Area of the ring = larger area - inner area
= pi(2.3^2-1.9^2) = 1.68(pi) = 5.278 sq cm
=======================================
Cheers,
Stan H.